# CrossIn: An Efficient Instruction Tuning Approach for Cross-Lingual Knowledge Alignment

### v1

@@ Abstract
Multilingual proficiency presents a significant challenge for large language models (LLMs). English-centric models are usually suboptimal in other languages, particularly those that are linguistically distant from English. This performance discrepancy mainly stems from the imbalanced distribution of training data across languages during pre-training and instruction tuning stages. To address this problem, we propose a novel approach called CrossIn, which utilizes a mixed composition of cross-lingual instruction tuning data. Our method leverages the compressed representation shared by various languages to efficiently enhance the model’s task-solving capabilities and multilingual proficiency within a single process. In addition, we introduce a multi-task and multi-faceted benchmark to evaluate the effectiveness of CrossIn. Experimental results demonstrate that our method substantially improves performance across tasks and languages, and we provide extensive insights into the impact of cross-lingual data volume and the integration of translation data on enhancing multilingual consistency and accuracy. [^1] The dataset and model information can be found at https://github.com/Lingy12/CrossIn.
CrossIn: An Efficient Instruction Tuning Approach for
Cross-Lingual Knowledge Alignment
@@ 1 Introduction
In this work, we propose an efficient approach based on a mixed composition of cross-lingual instruction tuning data to exploit LLMs’ underlying multilingual capability, which particularly improves the cross-lingual knowledge alignment qi2023crosslingual; wang2023seaeval. Instruction tuning is to boost the task solving capability of pre-trained language backbones alpaca; luo2023empirical; touvron2023llama. The task and prompt diversity are crucial in both data preparation and the training process, and a small high-quality set is sufficient to achieve state-of-the-art zero-shot performance ouyang2022training. However, the language diversity of instruction tuning is often overlooked in English-centric LLMs. We thus aim to enrich instruction tuning from the language perspective. Since all languages share the compressed representation space, cross-lingual instruction tuning can efficiently boost the model’s task-solving and multilingual capabilities within a single process. Unlike previous work that involved a multi-task setting by adding machine translation and mixing monolingual samples of each language zhu2023extrapolating, we integrate two languages at the sample level and combine various languages at the corpus level. Moreover, we compare various mixing strategies to identify the impact of different data formulations.
@@
  • Mixed Cross-Lingual Instruction Tuning. We introduce CrossIn, a cross-lingual instruction tuning approach aimed at aligning knowledge across languages to stimulate the model’s full multilingual capability after pre-training. It offers a more efficient way of improving the model’s performance in various linguistic contexts.
@@ 2.1 Multilingual Large Language Model
Instruction tuning on English datasets can introduce zero-shot capabilities in other languages as well wei2022finetuned; chung2022scaling. Further studies have explored the use of diverse training sets in multiple languages can improve cross-lingual generalization, suggesting that incorporating data from various languages can significantly enhance the model’s ability to generalize across linguistic boundaries muennighoff2023crosslingual; kew2023turning; shaham2024multilingual. In our work, we build upon these findings and focus on improving multilingual consistency through targeted instruction finetuning. By refining the instruction processing mechanism, we aim to enforce the alignment across different languages to improve multilingual capabilities.

### v2

@@ Abstract
Multilingual proficiency in large language models (LLMs) presents a significant challenge due to the uneven distribution of training data and the English-centric focus during instruction tuning. To mitigate these issues, we introduce CrossIn (Cross-lingual Instruction Tuning), which utilizes two distinct types instruction tuning datasets: the Complex Task Dataset (CTD), rich in diverse, high-quality and logical tasks like math and coding, and the Linguistic Uniformity Dataset (LUD), consisting of easier-to-translate, linguistically uniform tasks. CrossIn merges cross-lingual instruction data from CTD with machine-translated data from LUD to reinforce knowledge alignment during instruction tuning. This strategy allows us to enhance reasoning capabilities without compromising knowledge consistency across languages. We also present a multi-task benchmark to evaluate CrossIn, with results showing substantial improvements in performance across languages and tasks. This demonstrates the benefits of integrating cross-lingual data and translation in enhancing multilingual consistency and accuracy during instruction tuning.  [^1] We plan to release our datasets and models following the conclusion of the anonymity period.
CrossIn: An Efficient Instruction Tuning Approach for Cross-Lingual Knowledge Alignment
@@ 1 Introduction
In this paper, we explore instruction tuning with two distinct datasets: the Complex Task Dataset (CTD), which contains a variety of high-quality, hard-to-translate tasks like math and coding, and the Linguistic Uniformity Dataset (LUD), characterized by easier-to-translate, linguistically uniform tasks. Tuning with CTD often leads to lower consistency due to model forgetting luo2024empirical, while tuning with LUD may lead to lower reasoning capability due to its homogeneous nature.
To address these issues, we propose a method that enhances both the consistency of cross-lingual instruction tuning and task performance accuracy. Our approach leverages advanced tuning strategies that exploit the logical structures within tasks, thereby improving logical reasoning and model effectiveness across different languages. This method not only balances the complexities associated with language-specific nuances but also enhances overall model performance in multilingual environments. Our results demonstrate substantial improvements in multilingual proficiency and task accuracy, advancing the capabilities of language models.
@@
  • Mixed Cross-Lingual Instruction Tuning. We introduce CrossIn, a cross-lingual instruction tuning approach aimed at aligning knowledge across languages to stimulate the model’s full multilingual capability after pre-training. It offers a more reliable way of improving the model’s capability without loss of consistency.
@@ 2.1 Multilingual Large Language Model
Instruction tuning using English datasets has demonstrated the potential to extend zero-shot capabilities across multiple languages wei2022finetuned; chung2022scaling, however, they do have some limitation which is discussed in LABEL:sec:data_5_2. Previous research supports the notion that utilizing training sets composed of diverse languages can significantly enhance cross-lingual generalization muennighoff2023crosslingual; kew2023turning; shaham2024multilingual. Building on these insights, our work focuses on enhancing multilingual consistency by specifically targeting instruction finetuning. By optimizing the instruction processing mechanism, we aim to ensure better alignment across different languages during instruction tuning phase.